### v4

@@ 1 Introduction
One widely used estimation technique for interactive fixed effect panel regressions is the least squares (LS) method,555 The LS estimator in this context is also sometimes called concentrated least squares estimator, and was originally proposed by Kiefer (1980). which treats $\lambda$ and $f$ as parameters to estimate (fixed effects). [^6] Other estimation methods of panel regressions with interactive fixed effects include the quasi-difference approach (e.g., Holtz-Eakin, Newey, and Rosen 1988), generalized method of moments estimation (e.g. Ahn, Lee, and Schmidt 2001, 2013), the common correlated random effect method (e.g., Pesaran 2006), the decision theoretic approach (e.g., Chamberlain and Moreira 2009), and Lasso type shrinkage methods on fixed effects (e.g., Cheng, Liao, and Schorfheide 2016, Lu and Su 2016, Su, Shi, and Phillips 2016). Let the Frobenius norm of an $N\times T$ matrix $A$ be $\|A\|_{2}:=\left(\sum_{i=1}^{N}\sum_{t=1}^{T}A_{it}^{2}\right)^{1/2}$. Then, the LS estimator for $\beta$ reads
@@
The LS estimation approach is convenient, because it does not restrict the relationship between the unobserved heterogeneity ($\Gamma_{0}$) and the observed explanatory variables ($X_{1},...,X_{K}$). However, the calculation of $\widehat{\beta}_{{\rm LS},R}$ requires solving a non-convex optimization problem. While $\left\|Y-\beta\cdot X-\Gamma\right\|_{2}^{2}$ is a convex function of $\beta$ and $\Gamma$, the profiled objective function $L_{R}(\beta)$ is in general not convex in $\beta$, and can have multiple local minima, as will be discussed in Section 2.1 in more detail. The reason for the non-convexity is that the constraint ${\rm rank}(\Gamma)\leq R$ is non-convex. This non-convexity can become a serious computational obstacle, for example, when the number of regressors is too large to allow for a simple grid-search over $\beta$ (for a given $\beta$, the optimization over $\Gamma$ is a principal components problem that can generally be solved quickly in the linear regression case). For generalizations to non-linear panel regression models (e.g., Chen 2014, Chen, Fernández-Val, and Weidner 2021), the optimization of the non-convex objective function with respect to the high-dimensional factors and loadings becomes even more challenging.
@@
Our first contribution is to overcome the non-convexity issue of the least squares estimation problem above by considering a panel regression with nuclear norm regularization, which provides a convex relaxation of the rank constraint in (4). To be more specific, let $s(\Gamma):=[s_{1}(\Gamma),\allowbreak s_{2}(\Gamma),\allowbreak\ldots,s_{\min(N,T)}(\Gamma)]$ be the vector of singular values of $\Gamma$777 The non-zero singular values of $\Gamma$ are the square roots of non-zero eigenvalues of $\Gamma\Gamma^{\prime}$. Singular values are non-negative by definition. The rank of a matrix is equal to the number of non-zero singular values, that is, ${\rm rank}(\Gamma)=\left\|s(\Gamma)\right\|_{0}$, where $\left\|v\right\|_{0}$ equals the number of non-zero elements of the vector $v$ (sometimes calles the “$\ell^{0}$-norm” of $v$). The nuclear norm of $\Gamma$ is defined by $\|\Gamma\|_{1}:=\left\|s(\Gamma)\right\|_{1}=\sum_{r=1}^{\min(N,T)}s_{r}(\Gamma)$, that is, the nuclear norm of the matrix $\Gamma$ is simply the $\ell^{1}$-norm of the vector $s(\Gamma)$.888 The nuclear norm $\|\Gamma\|_{1}$ is the convex envelope of ${\rm rank}(\Gamma)$ over the set of matrices with spectral norm at most one, see e.g. Recht, Fazel, and Parrilo (2010). The nuclear norm is also sometimes called trace norm, Schatten 1-norm, or Ky Fan n-norm. Our index notation for the nuclear norm $\|\Gamma\|_{1}$, Frobenius norm $\|\Gamma\|_{2}$, and spectral norm $\|\Gamma\|_{\infty}=\lim_{q\rightarrow\infty}\|\Gamma\|_{q}$ is motivated by the unifying formula $\|\Gamma\|_{q}=\left\{\sum_{r=1}^{\min(N,T)}\left[s_{r}(\Gamma)\right]^{q}\right\}^{1/q}$. A convex relaxation of (4) can then be obtained by replacing the non-convex constraint ${\rm rank}(\Gamma)\leq R$ by the convex constraint $\|\Gamma\|_{1}\leq c$, for some constant $c$. This gives
@@
We also define $\widehat{\beta}_{*}:=\lim_{\psi\rightarrow 0}\widehat{\beta}_{\psi}$ for fixed $N$ and $T$101010 Here, the limit $\psi\rightarrow 0$ is for fixed $N$ and $T$, and has nothing to do with our large $N$, $T$ asymptotic considerations. We will show in Section 2.2 that $\widehat{\beta}_{*}=\operatorname*{argmin}_{\beta}\left\|Y-\beta\cdot X\right\|_{1}$, that is, $\widehat{\beta}_{*}$ can alternatively be obtained by minimizing the nuclear norm of $Y-\beta\cdot X$. The nuclear norm minimizing estimator $\widehat{\beta}_{*}$ is novel to the literature as far as we know.
@@
Nuclear norm penalized estimation has been widely studied in the machine learning and statistical learning literature. There, the parameter of interest is usually the matrix that we call $\Gamma$ in our model. In particular, there are many papers that use this penalization method in matrix completion (e.g., Recht, Fazel, and Parrilo 2010 and Hastie, Tibshirani, and Wainwright 2015 for recent surveys), and for reduced rank regression estimation (e.g., Rohde and Tsybakov 2011). More recently, nuclear norm penalization has also been used in the econometrics literature: Bai and Ng (2017, 2019) use it to improve estimation in a pure factor model. Athey, Bayati, Doudchenko, Imbens, and Khosravi (2021) apply nuclear norm penalization to treatment effect estimation with unbalanced panel data due to missing observations together with a regularization on the high dimensional regression coefficients – their primary interest is to predict the left-hand side variable using the regularization. Chernozhukov, Hansen, Liao, and Zhu (2018) consider panel regression models with heterogeneous coefficients, while in this paper we focus on panel regression with homogenous coefficients. To the best of our knowledge, our results here on the estimates of the common regression coefficients $\beta_{0}$ are new in this literature, and the nuclear norm minimizing estimator $\widehat{\beta}_{*}$ has also not been proposed previously.
@@
Since the 2018 working paper version of this paper, many related studies have used nuclear norm regularization and similar methods in panel data, factor models, and network models. In network settings, Alidaee, Auerbach, and Leung (2020) and Ma, Su, and Zhang (2022) use penalized methods to recover structure in networks. In panel data models, Chernozhukov, Hansen, Liao, and Zhu (2018) use low-rank methods to study models with varying slopes, and Belloni, Chen, Madrid Padilla, and Wang (2023), Wang, Su, and Zhang (2022), and Feng (2024) extend this to quantile regressions, which relates to our work in Section 6. Miao, Li, and Su (2020) and Miao, Phillips, and Su (2023) look at threshold models and VARs with factor structures. Chen (2022) gives a general framework for conditional factor models using nuclear norm regularization, and also shows how to choose tuning parameters and prove consistency. Beyhum and Gautier (2022) allow for weak (non-strong) factors, and Fernández-Val, Freeman, and Weidner (2021) use nuclear norm penalties to estimate nonseparable panel models. Beyhum and Gautier (2019) treat the low-rank structure as an approximation, allowing the rank to grow with sample size. Chetverikov and Manresa (2022) and Mugnier (2022) study grouped fixed effects, and use nuclear norm methods in the first step of their slope estimation. Armstrong, Weidner, and Zeleneev (2022) also use the nuclear norm to make slope estimates more robust in interactive fixed effects models. Hong, Su, and Jiang (2023) develop a profile GMM method that uses nuclear norm regularization and allows for endogeneity. Vogt, Walsh, and Linton (2022) build a high-dimensional version of the CCE estimator, combining factor projections with Lasso. Gao, Peng, and Yan (2024) study robust inference when the errors are not Gaussian and may be dependent over time and across units, using thresholded long-run covariance estimation. Finally, Chen (2022) (already mentioned above) develop a flexible method that applies to many types of conditional factor models. Overall, this recent work shows how nuclear norm regularization and related convex methods are useful for many problems involving panel data, factor models, and networks.
@@
The paper is organized as follows. Section 2 discusses theoretical motivations for employing nuclear norm regularization instead conventional rank restrictions. In Section 3 we derive consistency results on $\widehat{\beta}_{\psi}$ and $\widehat{\beta}_{*}$. Section 4 shows how to use these two estimates as a preliminary step to construct an estimator through iterations that achieves asymptotic equivalence to the fixed effect estimator. Section 5 investigates finite sample properties via simulations. We then briefly discuss extensions to non-linear panel models with interactive fixed effects in Section 6, and Section 7 concludes the paper. All technical derivations and proofs are presented in the appendix or supplementary appendix.
@@ 2.1 Convex Relaxation
that is, the nuclear norm is dual to the spectral norm $\|\cdot\|_{\infty}$. From this it is easy to see that $\|\cdot\|_{1}$ is indeed a matrix norm, and thus convex in $\Gamma$111111 Let $B$ and $C$ be matrices of the same size. Then, by (6) there exists a matrix $A$ of the same size with $\|A\|_{\infty}\leq 1$ such that $\|B+C\|_{1}={\rm Tr}[A^{\prime}\,(B+C)]={\rm Tr}(A^{\prime}B)+{\rm Tr}(A^{\prime}C)\leq\|B\|_{1}+\|C\|_{1}$, which is the triangle inequality for the nuclear norm. Together with absolute homogeneity of $\left\|\cdot\right\|_{1}$ this implies convexity. Therefore, the nuclear norm regularized objective function
@@
where the largest $R$ singular values are omitted in the sum - because they are absorbed by the principal component estimates $\widehat{\lambda}$ and $\widehat{f}$. The remaining problem in calculating $\widehat{\beta}_{{\rm LS},R}$ is the generally non-convex minimization of $L_{R}(\beta)$ over $\beta$131313 In our discussion here we focus on the calculation of $\widehat{\beta}_{{\rm LS},R}$ via minimization of the profile objective function $L_{R}(\beta)$. More generally, $\widehat{\beta}_{{\rm LS},R}$ can be obtained by any method that minimizes $\left\|Y-\beta\cdot X-\lambda f^{\prime}\right\|_{2}^{2}$ over $\beta$, $\lambda$, $f$, see e.g. Bai (2009) or the supplementary appendix in Moon and Weidner (2015). For any such method the non-convexity of the objective function is a potential problem, because the algorithm may converge to a local minimum, or potentially even to a critical point that is not a local minimum. To illustrate the potential difficulty caused by this non-convexity, in Figure 1 we plot $L_{R}(\beta)$ for the simple example described in Appendix A.1. In this example $L_{R}(\beta)$ is non-convex and has two local minima, one of which (the global one) is close to the true parameter $\beta_{0}=2$. The figure also shows that $Q_{\psi}(\beta)$ is convex and only has a single local minimum.
@@
$\ell_{\psi}(s)$ :=$:=\left\{\begin{array}[]{ll}\frac{1}{2}\;s^{2},&\text{for $s<\psi$,}\\ 0,&\text{for $s\geq\psi$,}\end{array}\right.$ $q_{\psi}(s):=\left\{\begin{array}[]{ll}\frac{1}{2}\;s^{2},&\text{for $s<\psi$,}\\ \psi s-\frac{\psi^{2}}{2},&\text{for $s\geq\psi$.}\end{array}\right.$ (12)
@@
The proof is given in the appendix. Figure 2 shows the functions $q_{\psi}(s)$ and $\ell_{\psi}(s)$ for real valued arguments $s$ and $\psi=1$. For values $s<\psi$ the functions are identical, but at $s=\psi$ the function $\ell_{\psi}(s)$ has a non-continuous jump, implying that $\ell_{\psi}(s)$ is non-convex, while $q_{\psi}(s)$ continues linearly for $s\geq\psi$, thus remaining convex.
@@ 2.2 Unique Matrix Separation
When estimating the interactive fixed effect model (1), in practice both $\beta_{0}$ and $R_{0}$ are unknown. Showing that $\beta_{0}$ and $R_{0}$ can be consistently estimated jointly is a difficult problem in general. [^15] The problem of joint identification of $\beta_{0}$ and $R_{0}$ is often avoided in the literature. Some papers (e.g. Bai 2009, Li, Qian, and Su 2016, Moon and Weidner 2017) assume that the number of factors $R_{0}$ is known when showing consistency for an estimator of $\beta_{0}$. Alternatively, Lu and Su (2016) allow for unknown $R_{0}$, but assume consistency of their estimator for $\beta_{0}$. Within the interactive fixed effects estimation framework this joint inference problem has only been successfully addressed when both of the following assumptions are satisfied:161616 Some existing estimation methods avoid specifying $R$ when estimating $\beta_{0}$, but always at the cost of some additional assumptions on the data generating process. For example, the common correlated effects estimator of Pesaran (2006) avoids choosing $R$, but requires assumptions on how the factors $f_{0}$ enter into the observed regressors $X_{k}$, and requires all regressors of interest to be high-rank.
@@
In the following we explain how the nuclear norm minimization approach overcomes the restrictions (C1), that is, how to estimate regression coefficients when $R_{0}$ is unknown. We already introduced $\widehat{\beta}_{*}=\lim_{\psi\rightarrow 0}\widehat{\beta}_{\psi}$ in Section 1. Using Lemma 1 we can now characterize $\widehat{\beta}_{*}$ differently. It is easy to see that $\lim_{\psi\rightarrow 0}\psi^{-1}q_{\psi}(s)=s$, for $s\in[0,\infty)$, and therefore $\lim_{\psi\rightarrow 0}\psi^{-1}q_{\psi}(A)=\|A\|_{1}$, for $A\in\mathbb{R}^{N\times T}$. Lemma 1 thus implies that $\lim_{\psi\rightarrow 0}\psi^{-1}Q_{\psi}(\beta)=\|(Y-\beta\cdot X)/\sqrt{NT}\|_{1}$. Another way to see this is as follows. According to (14), the limit $\psi\rightarrow 0$ corresponds to choosing $R$ very large, i.e., $R=\min(N,T)$. In this case, $\widehat{\Gamma}_{\psi}(\beta)=Y-\beta\cdot X$, and the profile objective function is $Q_{\psi}(\beta)=\frac{\psi}{\sqrt{NT}}\|\widehat{\Gamma}_{\psi}(\beta)\|_{1}=\frac{\psi}{\sqrt{NT}}\|Y-\beta\cdot X\|_{1}.$ From this we deduce $\lim_{\psi\rightarrow 0}\psi^{-1}Q_{\psi}(\beta)=\frac{\|Y-\beta\cdot X\|_{1}}{\sqrt{NT}}.$
@@
We cannot expect the LS estimator $\widehat{\beta}_{{\rm LS},R}$ to have good properties (in particular consistency) if we choose the number of factors equal to, or close to, its maximum possible value $R=\min(N,T)$. It is therefore somewhat surprising that $\widehat{\beta}_{\psi}$ has a well-defined limit as $\psi\rightarrow 0$, and that we are able to show consistency of the limiting estimator $\widehat{\beta}_{*}$ under appropriate regularity conditions in the following sections, because the resulting estimator for $\Gamma$ is certainly not consistent for $\Gamma_{0}$ in that limit.181818 The $\psi\rightarrow 0$ limit (for fixed $N$, $T$) of the optimal $\Gamma$ in (5) is $Y-\widehat{\beta}_{*}\cdot X$, which as $N$ and $T$ grow converges to $\lambda_{0}f_{0}^{\prime}+E$ for consistent $\widehat{\beta}_{*}$, that is, the estimator for $\Gamma$ that corresponds to $\widehat{\beta}_{*}$ is not consistent for $\lambda_{0}f_{0}^{\prime}$.
@@
Next, we explain how the nuclear norm regularization approach helps to overcome the restrictions (C2) above, that is, how to estimate regression coefficients for low-rank regressors when $R_{0}$ is unknown. The goal is to provide conditions on the regressors $X_{k}$ under which the nuclear norm penalization approach indeed solves the matrix separation problem for low-rank regressors and interactive fixed effects.
We first want to answer this in a simplified setting, where the objective function is replaced by the expected objective function, that is, we consider

### v5

@@ 1 Introduction
One widely used estimation technique for interactive fixed effect panel regressions is the least squares (LS) method, [^5] The LS estimator in this context is also sometimes called concentrated least squares estimator, and was originally proposed by Kiefer (1980). which treats $\lambda$ and $f$ as parameters to estimate (fixed effects). [^6] Other estimation methods of panel regressions with interactive fixed effects include the quasi-difference approach (e.g., Holtz-Eakin, Newey, and Rosen 1988), generalized method of moments estimation (e.g. Ahn, Lee, and Schmidt 2001, 2013), the common correlated random effect method (e.g., Pesaran 2006), the decision theoretic approach (e.g., Chamberlain and Moreira 2009), and Lasso type shrinkage methods on fixed effects (e.g., Cheng, Liao, and Schorfheide 2016, Lu and Su 2016, Su, Shi, and Phillips 2016). Let the Frobenius norm of an $N\times T$ matrix $A$ be $\|A\|_{2}:=\left(\sum_{i=1}^{N}\sum_{t=1}^{T}A_{it}^{2}\right)^{1/2}$. Then, the LS estimator for $\beta$ reads
@@
The LS estimation approach is convenient, because it does not restrict the relationship between the unobserved heterogeneity ($\Gamma_{0}$) and the observed explanatory variables ($X_{1},...,X_{K}$). However, the calculation of $\widehat{\beta}_{{\rm LS},R}$ requires solving a non-convex optimization problem. While $\left\|Y-\beta\cdot X-\Gamma\right\|_{2}^{2}$ is a convex function of $\beta$ and $\Gamma$, the profiled objective function $L_{R}(\beta)$ is in general not convex in $\beta$, and can have multiple local minima, as will be discussed in Section 2.1 in more detail. The reason for the non-convexity is that the constraint ${\rm rank}(\Gamma)\leq R$ is non-convex. In practice, researchers often use iterative methods that alternate between estimating $\beta$ given $\Gamma$ and updating $\Gamma$ given $\beta$. However, due to the non-convexity, such iterative procedures may converge to local minima or inconsistent estimators if initialized with arbitrary values. The nuclear norm regularization approach proposed in this paper provides a consistent initial estimator, which can then be used to initialize iterative procedures that are guaranteed to converge to the correct solution. This non-convexity can also become a serious computational obstacle, for example, when the number of regressors is too large to allow for a simple grid-search over $\beta$ (for a given $\beta$, the optimization over $\Gamma$ is a principal components problem that can generally be solved quickly in the linear regression case). For generalizations to non-linear panel regression models (e.g., Chen 2014, Chen, Fernández-Val, and Weidner 2021), the optimization of the non-convex objective function with respect to the high-dimensional factors and loadings becomes even more challenging.
@@
Our first contribution is to overcome the non-convexity issue of the least squares estimation problem above by considering a panel regression with nuclear norm regularization, which provides a convex relaxation of the rank constraint in (4). To be more specific, let $s(\Gamma):=[s_{1}(\Gamma),\allowbreak s_{2}(\Gamma),\allowbreak\ldots,s_{\min(N,T)}(\Gamma)]$ be the vector of singular values of $\Gamma$. [^7] The non-zero singular values of $\Gamma$ are the square roots of non-zero eigenvalues of $\Gamma\Gamma^{\prime}$. Singular values are non-negative by definition. The rank of a matrix is equal to the number of non-zero singular values, that is, ${\rm rank}(\Gamma)=\left\|s(\Gamma)\right\|_{0}$, where $\left\|v\right\|_{0}$ equals the number of non-zero elements of the vector $v$ (sometimes calles the $\ell^{0}$-norm” of $v$). The nuclear norm of $\Gamma$ is defined by $\|\Gamma\|_{1}:=\left\|s(\Gamma)\right\|_{1}=\sum_{r=1}^{\min(N,T)}s_{r}(\Gamma)$, that is, the nuclear norm of the matrix $\Gamma$ is simply the $\ell^{1}$-norm of the vector $s(\Gamma)$. [^8] The nuclear norm $\|\Gamma\|_{1}$ is the convex envelope of ${\rm rank}(\Gamma)$ over the set of matrices with spectral norm at most one, see e.g. Recht, Fazel, and Parrilo (2010). The nuclear norm is also sometimes called trace norm, Schatten 1-norm, or Ky Fan n-norm. Our index notation for the nuclear norm $\|\Gamma\|_{1}$, Frobenius norm $\|\Gamma\|_{2}$, and spectral norm $\|\Gamma\|_{\infty}=\lim_{q\rightarrow\infty}\|\Gamma\|_{q}$ is motivated by the unifying formula $\|\Gamma\|_{q}=\left\{\sum_{r=1}^{\min(N,T)}\left[s_{r}(\Gamma)\right]^{q}\right\}^{1/q}$. A convex relaxation of (4) can then be obtained by replacing the non-convex constraint ${\rm rank}(\Gamma)\leq R$ by the convex constraint $\|\Gamma\|_{1}\leq c$, for some constant $c$. This gives
@@
We also define $\widehat{\beta}_{*}:=\lim_{\psi\rightarrow 0}\widehat{\beta}_{\psi}$ for fixed $N$ and TT. [^10] Here, the limit $\psi\rightarrow 0$ is for fixed $N$ and $T$, and has nothing to do with our large $N$, $T$ asymptotic considerations. We will show in Section 2.2 that $\widehat{\beta}_{*}=\operatorname*{argmin}_{\beta}\left\|Y-\beta\cdot X\right\|_{1}$, that is, $\widehat{\beta}_{*}$ can alternatively be obtained by minimizing the nuclear norm of $Y-\beta\cdot X$. The nuclear norm minimizing estimator $\widehat{\beta}_{*}$ is novel to the literature as far as we know.
@@
Nuclear norm penalized estimation has been widely studied in the machine learning and statistical learning literature. There, the parameter of interest is usually the matrix that we call $\Gamma$ in our model. In particular, there are many papers that use this penalization method in matrix completion (e.g., Recht, Fazel, and Parrilo 2010 and Hastie, Tibshirani, and Wainwright 2015 for recent surveys), and for reduced rank regression estimation (e.g., Rohde and Tsybakov 2011). More recently, nuclear norm penalization has also been used in the econometrics literature: Bai and Ng (2019) use it to improve estimation in a pure factor model. Athey, Bayati, Doudchenko, Imbens, and Khosravi (2021) apply nuclear norm penalization to treatment effect estimation with unbalanced panel data due to missing observations together with a regularization on the high dimensional regression coefficients – their primary interest is to predict the left-hand side variable using the regularization. Chernozhukov, Hansen, Liao, and Zhu (2023) consider panel regression models with heterogeneous coefficients, while in this paper we focus on panel regression with homogenous coefficients. To the best of our knowledge, our results here on the estimates of the common regression coefficients $\beta_{0}$ are new in this literature, and the nuclear norm minimizing estimator $\widehat{\beta}_{*}$ has also not been proposed previously.
@@
Since the 2018 working paper version of this paper, many related studies have used nuclear norm regularization and similar methods in panel data, factor models, and network models. In network settings, Alidaee, Auerbach, and Leung (2020) and Ma, Su, and Zhang (2022) use penalized methods to recover structure in networks. In panel data models, Chernozhukov, Hansen, Liao, and Zhu (2023) use low-rank methods to study models with varying slopes, and Belloni, Chen, Madrid Padilla, and Wang (2023), Wang, Su, and Zhang (2022), and Feng (2024) extend this to quantile regressions, which relates to our work in Appendix B. Miao, Li, and Su (2020) and Miao, Phillips, and Su (2023) look at threshold models and VARs with factor structures. Chen (2022) gives a general framework for conditional factor models using nuclear norm regularization, and also shows how to choose tuning parameters and prove consistency. Beyhum and Gautier (2022) allow for weak (non-strong) factors, and Fernández-Val, Freeman, and Weidner (2021) use nuclear norm penalties to estimate nonseparable panel models. Beyhum and Gautier (2019) treat the low-rank structure as an approximation, allowing the rank to grow with sample size. Chetverikov and Manresa (2022) and Mugnier (2025) study grouped fixed effects, and use nuclear norm methods in the first step of their slope estimation. Armstrong, Weidner, and Zeleneev (2022) also use the nuclear norm to make slope estimates more robust in interactive fixed effects models. Hong, Su, and Jiang (2023) develop a profile GMM method that uses nuclear norm regularization and allows for endogeneity. Vogt, Walsh, and Linton (2022) build a high-dimensional version of the CCE estimator, combining factor projections with Lasso. Gao, Peng, and Yan (2024) study robust inference when the errors are not Gaussian and may be dependent over time and across units, using thresholded long-run covariance estimation. Finally, Chen (2022) (already mentioned above) develop a flexible method that applies to many types of conditional factor models. Overall, this recent work shows how nuclear norm regularization and related convex methods are useful for many problems involving panel data, factor models, and networks.
@@
The paper is organized as follows. Section 2 discusses theoretical motivations for employing nuclear norm regularization instead conventional rank restrictions. In Section 3 we derive consistency results on $\widehat{\beta}_{\psi}$ and $\widehat{\beta}_{*}$. Section 4 shows how to use these two estimates as a preliminary step to construct an estimator through iterations that achieves asymptotic equivalence to the fixed effect estimator. Section 5 investigates finite sample properties via simulations, and Section 6 concludes the paper. All technical derivations and proofs are presented in the appendix or supplementary appendix.
@@ 2.1 Convex Relaxation
that is, the nuclear norm is dual to the spectral norm $\|\cdot\|_{\infty}$. From this it is easy to see that $\|\cdot\|_{1}$ is indeed a matrix norm, and thus convex in $\Gamma$. [^11] Let $B$ and $C$ be matrices of the same size. Then, by (6) there exists a matrix $A$ of the same size with $\|A\|_{\infty}\leq 1$ such that $\|B+C\|_{1}={\rm Tr}[A^{\prime}\,(B+C)]={\rm Tr}(A^{\prime}B)+{\rm Tr}(A^{\prime}C)\leq\|B\|_{1}+\|C\|_{1}$, which is the triangle inequality for the nuclear norm. Together with absolute homogeneity of $\left\|\cdot\right\|_{1}$ this implies convexity. Therefore, the nuclear norm regularized objective function
@@
where the largest $R$ singular values are omitted in the sum - because they are absorbed by the principal component estimates $\widehat{\lambda}$ and $\widehat{f}$. The remaining problem in calculating $\widehat{\beta}_{{\rm LS},R}$ is the generally non-convex minimization of $L_{R}(\beta)$ over $\beta$. [^13] In our discussion here we focus on the calculation of $\widehat{\beta}_{{\rm LS},R}$ via minimization of the profile objective function $L_{R}(\beta)$. More generally, $\widehat{\beta}_{{\rm LS},R}$ can be obtained by any method that minimizes $\left\|Y-\beta\cdot X-\lambda f^{\prime}\right\|_{2}^{2}$ over $\beta$, $\lambda$, $f$, see e.g. Bai (2009) or the supplementary appendix in Moon and Weidner (2015). For any such method the non-convexity of the objective function is a potential problem, because the algorithm may converge to a local minimum, or potentially even to a critical point that is not a local minimum. To illustrate the potential difficulty caused by this non-convexity, in Figure 1 we plot $L_{R}(\beta)$ for the simple example described in Appendix A.1. In this example $L_{R}(\beta)$ is non-convex and has two local minima, one of which (the global one) is close to the true parameter $\beta_{0}=2$. The figure also shows that $Q_{\psi}(\beta)$ is convex and only has a single local minimum.
@@
$\ell_{\psi}(s)$ $:=\left\{\begin{array}[]{ll}\frac{1}{2}\;s^{2},&\text{for $s<\psi$,}\\ 0,&\text{for $s\geq\psi$,}\end{array}\right$. $q_{\psi}(s):=\left\{\begin{array}[]{ll}\frac{1}{2}\;s^{2},&\text{for $s<\psi$,}\\ \psi s-\frac{\psi^{2}}{2},&\text{for $s\geq\psi$.}\end{array}\right$. (12)
@@
The proof is given in the appendix. We note that Lemma 1 implies convexity of $Q_{\psi}(\beta)$ in $\beta$, since $(Y-\beta\cdot X)/\sqrt{NT}$ is linear in $\beta$ and $A\mapsto\sum_{r}q_{\psi}(s_{r}(A))$ is convex in $A$. The latter follows from a standard result in convex analysis: if $g:\mathbb{R}\rightarrow\mathbb{R}$ is convex and nondecreasing on $0,\infty)$, then $A\mapsto\sum_{r}g(s_{r}(A))$ is convex in the matrix $A$ (see, e.g., Lewis 1995). Figure 2 shows the functions $q_{\psi}(s)$ and $\ell_{\psi}(s)$ for real valued arguments $s$ and $\psi=1$.
@@ 2.2 Unique Matrix Separation
When estimating the interactive fixed effect model (1), in practice both $\beta_{0}$ and $R_{0}$ are unknown. Showing that $\beta_{0}$ and $R_{0}$ can be consistently estimated jointly is a difficult problem in general. [^15] The problem of joint identification of $\beta_{0}$ and $R_{0}$ is often avoided in the literature. Some papers (e.g. Bai 2009, Li, Qian, and Su 2016, Moon and Weidner 2017) assume that the number of factors $R_{0}$ is known when showing consistency for an estimator of $\beta_{0}$. Alternatively, Lu and Su (2016) allow for unknown $R_{0}$, but assume consistency of their estimator for $\beta_{0}$. Within the interactive fixed effects estimation framework this joint inference problem has only been successfully addressed when both of the following assumptions are satisfied: [^16] Some existing estimation methods avoid specifying $R$ when estimating $\beta_{0}$, but always at the cost of some additional assumptions on the data generating process. For example, the common correlated effects estimator of Pesaran (2006) avoids choosing $R$, but requires assumptions on how the factors $f_{0}$ enter into the observed regressors $X_{k}$, and requires all regressors of interest to be high-rank.
@@
Both the nuclear norm regularized estimator $\widehat{\beta}_{\psi}$ and the nuclear norm minimizing estimator $\widehat{\beta}_{*}$ can be computed without specifying the number of factors $R_{0}$, thus avoiding the restriction (C1). In this subsection we characterize $\widehat{\beta}_{*}$ and provide intuition for why these estimators can achieve consistency even when $R_{0}$ is unknown. We already introduced $\widehat{\beta}_{*}=\lim_{\psi\rightarrow 0}\widehat{\beta}_{\psi}$ in Section 1. Using Lemma 1 we can now characterize $\widehat{\beta}_{*}$ differently. It is easy to see that $\lim_{\psi\rightarrow 0}\psi^{-1}q_{\psi}(s)=s$, for $s\in[0,\infty)$, and therefore $\lim_{\psi\rightarrow 0}\psi^{-1}q_{\psi}(A)=\|A\|_{1}$, for $A\in\mathbb{R}^{N\times T}$. Lemma 1 thus implies that $\lim_{\psi\rightarrow 0}\psi^{-1}Q_{\psi}(\beta)=\|(Y-\beta\cdot X)/\sqrt{NT}\|_{1}$. Another way to see this is as follows. According to (14), the limit $\psi\rightarrow 0$ corresponds to choosing $R$ very large, i.e., $R=\min(N,T)$. In this case, $\widehat{\Gamma}_{\psi}(\beta)=Y-\beta\cdot X$, and the profile objective function is $Q_{\psi}(\beta)=\frac{\psi}{\sqrt{NT}}\|\widehat{\Gamma}_{\psi}(\beta)\|_{1}=\frac{\psi}{\sqrt{NT}}\|Y-\beta\cdot X\|_{1}.$ From this we deduce $\lim_{\psi\rightarrow 0}\psi^{-1}Q_{\psi}(\beta)=\frac{\|Y-\beta\cdot X\|_{1}}{\sqrt{NT}}.$
@@
We cannot expect the LS estimator $\widehat{\beta}_{{\rm LS},R}$ to have good properties (in particular consistency) if we choose the number of factors equal to, or close to, its maximum possible value $R=\min(N,T)$. It is therefore somewhat surprising that $\widehat{\beta}_{\psi}$ has a well-defined limit as $\psi\rightarrow 0$, and that we are able to show consistency of the limiting estimator $\widehat{\beta}_{*}$ under appropriate regularity conditions in the following sections, because the resulting estimator for $\Gamma$ is certainly not consistent for $\Gamma_{0}$ in that limit. [^18] The $\psi\rightarrow 0$ limit (for fixed $N$, $T$) of the optimal $\Gamma$ in (5) is $Y-\widehat{\beta}_{*}\cdot X$, which as $N$ and $T$ grow converges to $\lambda_{0}f_{0}^{\prime}+E$ for consistent $\widehat{\beta}_{*}$, that is, the estimator for $\Gamma$ that corresponds to $\widehat{\beta}_{*}$ is not consistent for $\lambda_{0}f_{0}^{\prime}$.
@@
To understand intuitively why $\widehat{\beta}_{*}$ can be consistent, consider that at any $\beta\neq\beta_{0}$ the residual matrix is $Y-\beta\cdot X=\Gamma_{0}+E+X\cdot(\beta_{0}-\beta)$. For $\widehat{\beta}_{*}$ to correctly identify $\beta_{0}$, we need that $\|Y-\beta\cdot X\|_{1}>\|Y-\beta_{0}\cdot X\|_{1}=\|\Gamma_{0}+E\|_{1}$ for all $\beta\neq\beta_{0}$, which requires that the term $X\cdot(\beta_{0}-\beta)$ does not cancel out singular values of $\Gamma_{0}+E$. Since $\Gamma_{0}$ is low-rank and $E$ is generically full-rank, this imposes different requirements depending on the rank of $X$: (i) if $X$ is low-rank, then $X$ must operate under sufficiently different row and column spaces than $\Gamma_{0}$, so that $X\cdot(\beta_{0}-\beta)$ cannot be absorbed into the factor structure; (ii) if $X$ is high-rank, then $X$ must be sufficiently different from $E$, so that $X\cdot(\beta_{0}-\beta)$ does not simply offset the idiosyncratic errors. The formal conditions in Section 3 make these requirements precise. Notice also that even when $\widehat{\beta}_{*}$ is consistent for $\beta_{0}$, the corresponding residual $Y-\widehat{\beta}_{*}\cdot X$ estimates $\Gamma_{0}+E$, not $\Gamma_{0}$ alone.
@@
Next, we provide an identification result showing how the nuclear norm regularization approach helps to overcome the restrictions (C2) above, that is, how to estimate regression coefficients for low-rank regressors when $R_{0}$ is unknown. The goal is to provide conditions on the regressors $X_{k}$ under which the nuclear norm penalization approach indeed solves the matrix separation problem for low-rank regressors and interactive fixed effects.
To build intuition, we first present an identification result based on the expected objective function for fixed $N$ and $T$. We consider